OpenStudy (anonymous):

find the limit of: lim x--> -4, n^2-16/sq root (n)-2

OpenStudy (turingtest):

\[\lim_{x\to-4}{n^2-16\over\sqrt n-2}\]correct?

OpenStudy (anonymous):

yes!

OpenStudy (turingtest):

try using difference of squares to factor the top

OpenStudy (anonymous):

(n-4) (n+4) what do I do on the bottom?

OpenStudy (turingtest):

nothing, factor the top with difference of squares again and watch the magic happen

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

please don't go.

OpenStudy (turingtest):

n-4 can be factored with difference of squares

OpenStudy (anonymous):

(n+2) (n-2), now what do I do.

OpenStudy (turingtest):

you didn't factor that quite right, try foiling it out and you'll see you don;t get n-4

OpenStudy (anonymous):

eek! I need hep on the factoring too.

OpenStudy (turingtest):

well difference of squares is\[a^2-b^2=(a+b)(a-b)\]what are a and b in your case of n-4 ?

OpenStudy (anonymous):

uuh n isn't squared, I do not know.4?

OpenStudy (turingtest):

compare\[a^2-b^2\]\[n-4\]so that means that \(b^2=4\) and \(a^2=n\), so then what are \(a\) and \(b\) ?

OpenStudy (anonymous):

b is 2, for 2 squared is 4. now, n, I do not know.

OpenStudy (turingtest):

\[a^2=n\implies a=\sqrt n\]make sense?

OpenStudy (anonymous):

so, hwo would I work out the problem? I mean, can you list thesteps?

OpenStudy (turingtest):

if you mean can I give you the complete solution that's a "no", sorry keep going from where we were...

OpenStudy (turingtest):

just looking at the numerator we had\[n^2-16=(n-4)(n+4)\]I just told you how to factor \((n-4)\), so do that and see what you have left over

OpenStudy (turingtest):

or did you not follow my explanation on how to factor that?

OpenStudy (anonymous):

fair enough, haha! okay, so ari have t(n-4) factored out to (sq rt (n) +4) and (sq rt (n) -4) ? is thacorrct?

OpenStudy (turingtest):

perfect! now looking at what was in the denominator, see what's going to happen?

OpenStudy (turingtest):

oh actually those should be 2's

OpenStudy (anonymous):

the (sq rt (n) -4) will b canceled out.

OpenStudy (turingtest):

\[n-4=(\sqrt n-2)(\sqrt n+2)\]

OpenStudy (anonymous):

leaving (sq rt (n) +4) and (n+4)

OpenStudy (turingtest):

those should be 2's, not 4's

OpenStudy (turingtest):

right?

OpenStudy (anonymous):

oops. okay.

OpenStudy (anonymous):

how, do I plug in the original -4?

OpenStudy (turingtest):

so yeah, the sqrt(n)-2 's cancel and what do we have left over?

OpenStudy (anonymous):

-4 -2?

OpenStudy (anonymous):

no.

OpenStudy (turingtest):

nope, let's go through it again :P

OpenStudy (anonymous):

0?

OpenStudy (turingtest):

yes the final answer is zero, you sure got that fast o.0 nice job though :) sure you understand?

OpenStudy (anonymous):

n + 4 and sq rt n +2 plugged in, correct?

OpenStudy (turingtest):

not sure what you mean... you have to plug in everything I think you got the right answer by a bit of luck; let's make sure you understand, okay?

OpenStudy (anonymous):

haha, ok.

OpenStudy (anonymous):

jlm dk mdiz'

OpenStudy (anonymous):

jopo d kf k w 'alfa

OpenStudy (turingtest):

\[\lim_{n\to-4}{n^2-16\over\sqrt n-2}=\lim_{n\to-4}{(n-4)(n+4)\over\sqrt n-2}=\lim_{n\to-4}{\cancel{(\sqrt n-2)}(\sqrt n+2)(n+4)\over\cancel{\sqrt n-2}}\]

OpenStudy (anonymous):

yes.

OpenStudy (turingtest):

oh I see what you meant, yes I guess you do just plug that in :P

OpenStudy (anonymous):

haha, okay thank you so much! you've great! nowi, I wont keep you! hel others the way you've helped me! :)

OpenStudy (turingtest):

I try, see ya around :)